OpenStudy (anonymous):

A rectangle that is 8 inches long and 5 inches wide has the same area as a triangle that is 2 inches high. What is the base of the triangle? Show and explain each step to find the answer.

OpenStudy (anonymous):

8x5=40 area of triangle = 1/2 bh h=2 40=1/2 2b 40=b The base of the triangle is 40 inches.

OpenStudy (anonymous):

thx